OpenStudy (quickstudent):

Can someone please help me with this problem? simplifying rational expressions using GCF

OpenStudy (quickstudent):

Can someone please walk me through this?

OpenStudy (quickstudent):

@TheSmartOne

TheSmartOne (thesmartone):

*

OpenStudy (quickstudent):

I found the GCF was x^3, this is what I got so far: \[\frac{ (x^3*x)+(x^3*x^3) }{ x^3 }\] But now I'm not sure which terms to cancel out.

OpenStudy (quickstudent):

@welshfella

OpenStudy (quickstudent):

@inkyvoyd

TheSmartOne (thesmartone):

Ok, now I'm back on OpenStudy again. :)

TheSmartOne (thesmartone):

\(\sf\Large \frac{x^4 + x^6}{x^3} = \frac{x^3(x+x^3)}{x^3} = \boxed{x +x^3}\)

TheSmartOne (thesmartone):

@quickstudent